undeadknight26 (undeadknight26):

A recursive function is shown below: f(1) = 3 and f(n) = f(n - 1) - 4; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 3, 1, 5, 9, 13, ... 3, 7, 11, 15, 19, ... 3, -7, -11, -15, -19, ... 3, -1, -5, -9, -13, ...

undeadknight26 (undeadknight26):

I think it is the last one.

OpenStudy (blackops2luvr):

no its c

OpenStudy (anonymous):

it's the last one

undeadknight26 (undeadknight26):

explanation please...and black ops 2 is boss :D

OpenStudy (blackops2luvr):

lol :D thnx

OpenStudy (anonymous):

f(2) = f(1) - 4 = 3 - 4 = -1

OpenStudy (blackops2luvr):

aww :I

OpenStudy (blackops2luvr):

he did it :D

OpenStudy (blackops2luvr):

medal?

undeadknight26 (undeadknight26):

1 more please? I think it is also the last one: Which equation is not equivalent to the formula d = rt? r equals t over d r equals d over t t equals d over r d = tr

OpenStudy (blackops2luvr):

i think its the last one too

undeadknight26 (undeadknight26):

ok thank you.

OpenStudy (blackops2luvr):

:3

OpenStudy (blackops2luvr):

\(\large\color{lime}\dagger\)\(\Large\color{navy}\dagger\)\(\huge\color{silver}\dagger\)\(\Huge\color{lime}\dagger\)\(\huge\color{silver}\dagger\)\(\Large\color{navy}\dagger\)\(\large\color{lime}\dagger\)